# Degeneracy Enhancement of Neutron-Antineutron Oscillation in Neutron Star

### v1

@@ Abstract
where $\Gamma\equiv 2Rn_{n}\langle\sigma v\rangle/(1+R)$ takes the role of neutron burning rate.
Suppose the neutron burning does not consume a significant fraction of the NS mass, the neutron density $n_{n}$ and the annihilation rate $\Gamma$ can then be treated as constants if the temperature is also a constant. Then the total nucleon number decays as, $N(t)=N(0)e^{-\Gamma t}$. With $n_{n}=10^{38}~{}\rm{cm^{-3}}$ and $\braket{\sigma v}\approx 10^{-15}~{}\rm{cm^{3}/s}$ Buccella:1987pv , the neutron burning rate can reach $\Gamma\approx 2R\times 10^{23}~{}{\rm{s^{-1}}}$ where we have taken the approximation $R\ll 1$. For NS to exist for more than one million years, $\Gamma\lesssim 3\times 10^{-14}$ s-1, which would require $R\lesssim 10^{-37}$. The actual constraint is even much stronger since the temperature descreases very fast with time and the resulting degeneracy parameter $(\mu-m)/T$ can reach above $10^{5}\sim 10^{6}$ as shown in Fig. 1.
@@
We implement the heating due to $n\bar{n}$ oscillation in NSCool Page:2004fy ; NSCool program to simulate the NS temperature evolution. As shown in Fig. 2, the heating effect by $n\bar{n}$ oscillation will prevent NS from cooling at some point. At the level of $10^{5}\sim 10^{6}$ years, the degeneracy parameter $(\mu-m)/T$ can reach $10^{5}\sim 10^{6}$. According to Eq. (11), the antineutron fraction $R_{0}$ is enhanced by a factor of $e^{2(\mu-m)/T}$ and hence its constraints can reach $\log_{10}R_{0}\sim-\mathcal{O}(10^{4})$ to be consistent with the observed NS cooling data Potekhin:2020ttj . When converted to the mixing angle, the constraint can reach $s^{2}\lesssim 10^{-\mathcal{O}(10^{4})}$.
Figure 2: The NS surface temperature data and cooling curves for both relativistic (solid) and non-relativistic (dashed) dispersion relations. Instead of the mixing angle $s^{2}$, the input parameter is taken as the antineutron density ratio $R_{0}$ with $T_{0}=9.9\times 10^{9}$ K at the beginning of the NS evolution and $n_{n}=5.8\times 10^{38}$ cm-3 around the NS center.
@@
Summary and Conclusions – Although the fermionic oscillation has been well established and experimental verified with neutrino oscillations, the previous study is only about the oscillation in vauum or with an effective matter potential at most. The description of $n\bar{n}$ oscillation is simply borrowed from its neutrino counterpart. However, most neutrons are confined inside nuclei or the compact NS with high degeneracy. It is then necessary for $n\bar{n}$ oscillation to consider the degenerate environment. Although our analysis takes the $n\bar{n}$ oscillation inside NS as illustration, the same formalism should also apply for nuclei.
Our result shows that this degeneracy effect can significantly change the neutron oscillation behavior. Especially, the probed GUT scale can go far beyond the Planck scale which means that those GUT models with $n\bar{n}$ oscillation are essentially excluded by the stability of NS. Otherwise, NS may explode when the degeneracy parameter $(\mu-m)/T$ significantly increases with much lower internal temperature towards the end of NS cooling curve. Similar conclusion should also apply for the neutron-mirror-neutron oscillation scenarios with cooling effect.
@@ Acknowledgements
The authors would like to thank Bo-Wen Ouyang, Pedro Pasquini, Jiaming Zheng, and Albert Zhou for useful discussions. The authors are supported by the National Natural Science Foundation of China (12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604). SFG is also an affiliate member of Kavli IPMU, University of Tokyo. XYF, ZYG, and QHW are supported by the Zhiyuan Future Scholar program (ZIRC2024-10) of the SJTU Zhiyuan College.

### v2

@@ Abstract
where $\Gamma\equiv 2Rn_{n}\langle\sigma v\rangle/(1+R)$ is the neutron burning rate.
Suppose the neutron burning does not consume a significant fraction of the NS mass, the neutron density $n_{n}$ and the annihilation rate $\Gamma$ can then be treated as constants if the temperature is also a constant. Then the total nucleon number decays as, $N(t)=N(0)e^{-\Gamma t}$. With $n_{n}=10^{38}\,\rm{cm^{-3}}$ and $\braket{\sigma v}\approx 10^{-15}\,\rm{cm^{3}/s}$ Buccella:1987pv , the neutron burning rate can reach $\Gamma\approx 2R\times 10^{23}\,{\rm{s^{-1}}}$ where we have taken the approximation $R\ll 1$. For NS to exist for more than one million years, $\Gamma\lesssim 3\times 10^{-14}$ s-1, which would require a lose constraint $R\lesssim 10^{-37}$. The actual constraint is even much stronger since the temperature descreases very fast with time and the resulting degeneracy parameter $(\mu-m)/T$ can reach above $10^{5}\sim 10^{6}$ Bertoni:2013bsa as shown in Fig. 2.
@@
Figure 3: The red-shifted NS surface temperature data and cooling curves for both relativistic (solid) and non-relativistic (dashed) dispersion relations. Instead of the mixing angle $s^{2}$, the input parameter is taken as the antineutron density ratio $R_{0}\equiv R(n_{0},T_{0})$ with $T_{0}=9.9\times 10^{9}$ K at the beginning of the NS evolution and $n_{0}=5.8\times 10^{38}$ cm-3 around the NS center.
We implement the heating due to $n\bar{n}$ oscillation in NSCool Page:2004fy ; NSCool program to simulate the NS temperature evolution. As shown in Fig. 3, the heating effect by $n\bar{n}$ oscillation will prevent NS from cooling at some point such that the cooling curve becomes flat. To be consistent with the observed data points Potekhin:2020ttj , the initial parameter $R_{0}$ should not exceed $\log_{10}R_{0}\sim-\mathcal{O}(10^{4})$. With $(\mu-m)/T$ reaching $10^{6}$ Bertoni:2013bsa around $10^{5}$ years, the physical antineutron fraction $R\approx(T/T_{0})^{3/2}e^{2(\mu-m)/T-2(\mu_{0}-m)/T_{0}}R_{0}$ is enhanced to $R\lesssim 10^{-43}$. Correspondingly, the heat release $2Rn^{2}_{n}\langle\sigma v\rangle Vm$ in the NS core with a radius of 2 km can reach $10^{24}$ W to far exceed the cooling power $4\pi R^{2}\sigma T^{4}\sim 10^{21}$ W at the NS surface. When converted to the mixing angle, the constraint can reach $s^{2}\lesssim 10^{-\mathcal{O}(10^{4})}$.
@@
Summary and Conclusions – We establish the formalism of degenerate oscillation in the QFT language with three different approaches of second quantization, thermal QFT, and kinetics theory. Although it can also apply for the bosonic case, we take the neutron-antineutron oscillation in NS for a concrete illustration. Our result shows that this degeneracy effect can significantly change the neutron oscillation behavior. Especially, the probed GUT scale can go far beyond the Planck scale which means that those GUT models with $n\bar{n}$ oscillation are essentially excluded by the stability of NS. Otherwise, NS may explode when the degeneracy parameter $(\mu-m)/T$ significantly increases with much lower internal temperature towards the end of NS cooling curve. Similar conclusion should also apply for the neutron-mirror-neutron oscillation scenarios with cooling effect.
@@ Acknowledgements
The authors would like to thank Bo-Wen Ouyang, Pedro Pasquini, Liang Tan, Jiaming Zheng, Albert Zhou, and Wei Zuo for useful discussions. SFG is particularly grateful to Luke Johns, Georg Raffelt, Guenter Sigl, and Maria Cristina Volpe for useful comments and discussions. The authors are supported by the National Natural Science Foundation of China (12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604). SFG is also an affiliate member of Kavli IPMU, University of Tokyo. XYF, ZYG, and QHW are supported by the Zhiyuan Future Scholar program (ZIRC2024-10) of the SJTU Zhiyuan College.